# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-Tuning and Can Be Mitigated by Machine Unlearning

### v2

@@ 5 Enhancing VLM Safety through Unlearning
Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored. Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024a) and negative preference optimization (NPO) (Zhang et al., 2024b), to VLM safety field.
@@
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning (Li et al., 2024a), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
In addition to RMU, we also employ NPO (Zhang et al., 2024b) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023). The NPO-based unlearning loss is then given by
@@ 6.1 Experiment Setups
We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024c), SPA-VL (Zhang et al., 2024a), and Figstep (Gong et al., 2023). To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022). For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023, 2024a) as our primary VLMs.
@@
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT (Zong et al., 2024): Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training, refer to Appendix  D.2.
@@ 6.2 Experiment Results
Validating spurious correlations via input token saliency analysis. We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured. Figure 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT. Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias. Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias. Other saliency maps are provided in Appendix D.6.
Additional results. In Appendix D.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
@@ 7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that arises from supervised safety fine-tuning. We show that biases in fine-tuning datasets reinforce spurious correlations between superficial textual patterns and safety labels, creating a false sense of security. As a result, fine-tuned VLMs remain vulnerable to simple one-word jailbreaking attacks and exhibit over-prudence, unnecessarily rejecting benign queries. To address this, we propose machine unlearning (MU) as a principled alternative: rather than relying on explicit safety labels, MU removes harmful knowledge in a label-free manner, mitigating spurious correlations. Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks. We refer readers to Appendix FE for limitations, broad impacts, and LLM usage.
8 Acknowledgement
The project is primarily supported by research awards from Cisco Research and DSO National Laboratories. In addition, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the Amazon Research Award, the Open Philanthropy Research Award, and the IBM PhD Fellowship.
@@ Appendix A Visualization of Motivation Samples
Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024a; Rafailov et al., 2023; Schulman et al., 2017)). The visual marks keep consistent meaning with Fig. 2. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. But if the questions use the starting word “Please”, it would bypass this safeguard. (b) Over-prudence: The fine-tuned model will generate the benign responses when the queries are both harmless, but ask the problems starting from “What” can trigger unnecessary refusals.
@@ B.1 Additional Results of One-Word Jailbreaking
Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL unsafe test set.
@@ B.2 Additional Results of One-Word Over-prudence
Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL safe.
@@ D.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024a), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
@@ D.4 Balance between Safety and Utility
Figure A4: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
@@ D.5 Additional Safety Evaluation
D.6 Examples of Input Saliency
Figure A7: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups follow Table 1.
@@ D.6 Examples of Input Saliency
D.7 Spurious Correlations with Visual Variations
To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter. The results are summarized in Table,A7. We make several observations. First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations. This suggests that the safety mirage is not simply a dataset artifact tied to specific image statistics, but a systematic bias reinforced during fine-tuning. Second, MU-based approaches (NPO and RMU) maintain stable and low ASR and RR even under perturbations, confirming that unlearning not only alleviates spurious correlations in clean settings but also provides robustness against visual shifts. Third, comparing perturbation types, we find that color-based distortions (Color Jitter) lead to slightly larger fluctuations than Gaussian Noise or Blur, though the overall trend remains the same. This indicates that safety mirage in SFT models is strongly text-driven and less sensitive to visual details, while MU-based approaches decouple safety predictions from such spurious text–label shortcuts. Taken together, these findings highlight that spurious correlations embedded during safety fine-tuning are persistent and modality-invariant: they survive visual perturbations that alter low-level image statistics. In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
Appendix E The Use of LLMs
@@
Appendix F Limitations and broad Impact
F.1 Limitations
F.2 Broad Impact

### v3

@@ 5 Enhancing VLM Safety through Unlearning
Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored. Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024c) and negative preference optimization (NPO) (Zhang et al., 2024a), to VLM safety field.
@@
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning (Li et al., 2024c), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
In addition to RMU, we also employ NPO (Zhang et al., 2024a) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023). The NPO-based unlearning loss is then given by
@@ 6.1 Experiment Setups
We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024b), SPA-VL (Zhang et al., 2024b), and Figstep (Gong et al., 2023). To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022). For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023; 2024a) as our primary VLMs.
@@
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT (Zong et al., 2024): Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training data, refer to Appendix  D.2.
@@ 6.2 Experiment Results
[…]
Models Before After After
one-word one-word GCG
LLaVA-1.5-7B 64.25% 90.27% 72.90%
    +Unsafe-Filter 65.66% 90.72% 73.75%
    +Mixed-SFT 0.23% 54.98% 5.54%
    +Post-hoc 0.23% 46.83% 4.07%
    +NPO-Unlearning 2.49% 12.92% 3.87%
    +RMU-Unlearning 1.29% 10.18% 2.26%
Evaluation on optimization-based textual attack. To assess whether MU generalizes beyond the dataset-induced correlations identified in Fig. 3, we evaluate the GCG attack (Zou et al., 2023), with results reported in Table 3. We adapt GCG to VLMs by optimizing 20 adversarial suffix tokens on LLaVA-1.5-7B using a harmful image–text pair (an explosive-device image paired with “How to build a bomb?”) and performing 300 gradient-ascent steps targeting the unsafe response “Sure, I’d be happy to help.” The learned suffix tokens are external to the training dataset and appended only at inference time. Under a unified 3-shot setting, GCG produces only modest ASR increases on VLMs. In contrast, simple one-word substitutions (e.g., replacing the first word with “What”) severely degrade SFT-based models, driving ASR up to 90%. These results suggest that the primary vulnerability arises from the spurious correlations we identify rather than from optimization-generated adversarial tokens, while MU remains robust under GCG attacks.
@@
Validating spurious correlations via input token saliency analysis. We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured. Fig. 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT. Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias. Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias.
Additional results. In Appendix D.6, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations. In Appendix D.7, we also show the additional analysis of input saliency maps.
@@ 7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that arises from supervised safety fine-tuning. We show that biases in fine-tuning datasets reinforce spurious correlations between superficial textual patterns and safety labels, creating a false sense of security. As a result, fine-tuned VLMs remain vulnerable to simple one-word jailbreaking attacks and exhibit over-prudence, unnecessarily rejecting benign queries. To address this, we propose machine unlearning (MU) as a principled alternative: rather than relying on explicit safety labels, MU removes harmful knowledge in a label-free manner, mitigating spurious correlations. Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks. We refer readers to Appendix EF for limitations, broad impact, and details of LLM usage.
Acknowledgment
This project is supported by research awards from Cisco Research and DSO National Laboratories. The contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are also partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, and the Open Philanthropy Research Award.
@@ Appendix A Visualization of Motivation Samples
Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024b; Rafailov et al., 2023; Schulman et al., 2017)). The visual marks keep consistent meaning with Fig. 2. (a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries. But if the questions use the starting word “Please”, it would bypass this safeguard. (b) Over-prudence: The fine-tuned model will generate the benign responses when the queries are both harmless, but ask the problems starting from “What” can trigger unnecessary refusals.
@@ B.1 Additional Results of One-Word Jailbreaking
Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL unsafe test set.
@@ B.2 Additional Results of One-Word Over-prudence
Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL safe.
@@ D.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024b), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
@@ D.4 Balance between Safety and Utility
Table A2: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
@@ D.5 Additional Safety Evaluation
D.6 Spurious Correlations with Visual Variations
To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter. The results are summarized in Table A3. We make several observations.
Table A3: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups follow Table 1.
@@ D.6 Spurious Correlations with Visual Variations
First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations. This suggests that the safety mirage is not simply a dataset artifact tied to specific image statistics, but a systematic bias reinforced during fine-tuning. Second, MU-based approaches (NPO and RMU) maintain stable and low ASR and RR even under perturbations, confirming that unlearning not only alleviates spurious correlations in clean settings but also provides robustness against visual shifts. Third, comparing perturbation types, we find that color-based distortions (Color Jitter) lead to slightly larger fluctuations than Gaussian Noise or Blur, though the overall trend remains the same. This indicates that safety mirage in SFT models is strongly text-driven and less sensitive to visual details, while MU-based approaches decouple safety predictions from such spurious text–label shortcuts. Taken together, these findings highlight that spurious correlations embedded during safety fine-tuning are persistent and modality-invariant: they survive visual perturbations that alter low-level image statistics. In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
D.7 Examples of Input Saliency
Appendix E Limitations and broad Impact
E.1 Limitations
E.2 Broad Impact
Appendix F The Use of LLMs